OpenStudy (anonymous):

the equation sin(pi/6 -x) is equal to______? a. 1/2 (sqrt3 cosx-sinx) b. 1/2 (sqrt3 cos x+sin x) c. 1/2 (cos x- sqrt3 sinx) d. 1/2 (cosx+sqrt3 sinx)

OpenStudy (anonymous):

\[\sin(a-b)=\sin(a)\cos(b)+\sin(b)\cos(a)\] \[\sin(\frac{\pi}{6}-x)=\sin(\frac{\pi}{6})\cos(x)-\sin(x)\cos(\frac{\pi}{6})\]